OpenStudy (babyslapmafro):

Please help me evaluate the following interval (click to see) using an appropriate substitution.

OpenStudy (babyslapmafro):

\[\int\limits_{}^{}\frac{ dx }{ \sqrt{x}e ^{(16\sqrt{x})} }\]

OpenStudy (babyslapmafro):

just realized i wrote interval and not integral

OpenStudy (anonymous):

\[u = 16\sqrt{x}\]\[du =\frac{ 8 dx }{ \sqrt{x} }\]\[8\int\limits_{}^{}\frac{ du }{ e ^{u} }\]

OpenStudy (anonymous):

So, just solve the new integral and resubstitute for the expression for x.

OpenStudy (babyslapmafro):

ok, and what does dx equal? Just so I know how to solve for dx...

OpenStudy (anonymous):

That's in the second of the three equations I determined.

OpenStudy (babyslapmafro):

Right, I do not know how to solve for dx

OpenStudy (anonymous):

You are in calculus, therefore, I'm sure you know basic arithmetic. Just rearrange the second equation. The three equations I wrote flow logically one to the next.

OpenStudy (babyslapmafro):

oh ok i see now... nvm

OpenStudy (anonymous):

hold on.

OpenStudy (anonymous):

My last equation is a little off, it should be\[\frac{ 1 }{ 8 }\int\limits_{}^{}\frac{ du }{ e ^{u} }\]

OpenStudy (babyslapmafro):

ok thanks

OpenStudy (anonymous):

uw!